# SystolicAttention: Fusing FlashAttention within a Single Systolic Array

### v2

@@ Abstract.
Transformer models rely heavily on scaled dot-product attention (SDPA), typically implemented using the FlashAttention algorithm. However, current systolic-array-based accelerators face significant challenges when executing FlashAttention. Systolic arrays can only achieve high utilization for consecutive and large matrix multiplications. In contrast, FlashAttention requires frequently interleaved matrix multiplications and softmax operations.
The frequent data swaps between the systolic array and external vector units result in low systolic array utilization. This is further exacerbated by the fact that softmax involves numerous non-matrix operations, which are not well-suited for systolic arrays. Moreover, the concurrent execution of matrix multiplication on systolic arrays and softmax on vector units leads to register file and SRAM port contention, further degrading performance.
To overcome these limitations, we propose FSA, an enhanced systolic array architecture that enables the entire FlashAttention algorithm to run entirely within a single systolic array—eliminating the need for external vector units. At the core of FSA is SystolicAttention, a novel scheduling algorithm that maps FlashAttention operations onto systolic arrays with fine-grained, element-wise overlap. This significantly improves array utilization while preserving the original floating-point operation order to maintain numerical stability.
We implement FSA in synthesizable RTL and evaluate its performance against state-of-the-art commercial accelerators. Our results show that FSA achieves 1.77$\times$ and 4.83$\times$ higher attention FLOPs/s utilization compared to AWS NeuronCore-v2 and Google’s TPUv5e, respectively, with only about 10% area overhead.
@@ 1. Introduction
The rapid advancement of artificial intelligence, particularly in natural language processing and computer vision, has been driven largely by the success of transformer models (Vaswani et al., 2017; Devlin et al., 2019; Brown et al., 2020). These models have revolutionized machine learning applications, from language translation and text generation (Ouyang et al., 2022) to image recognition (Kolesnikov et al., 2021) and autonomous driving (Carion et al., 2020). However, the computational demands of transformer models are enormous (Strubell et al., 2019), requiring specialized hardware accelerators to achieve practical performance and energy efficiency. As these models continue to grow in size and complexity (Kaplan et al., 2020), the need for efficient hardware acceleration becomes increasingly critical.
@@
Systolic arrays (Kung, 1982) are specialized hardware architectures that are designed to efficiently perform matrix multiplication by streaming data through a network of processing elements (PEs). In these architectures, data moves only one hop per clock cycle, minimizing hardware data movement overhead and enabling high scalability. Due to their efficiency and scalability, large systolic arrays have been widely adopted in AI accelerators. Notable examples include Google’s TPU (Google Cloud, 2024; Jouppi et al., 2017, 2023) (256$\times$256 in TPUv6e and 128$\times$128 prior to v6e) and AWS NeuronCore (AWS Neuron Documentation, 2024b) (128$\times$128).
@@
One of the key operations in transformer models is the scaled dot-product attention (SDPA), which is often implemented using the FlashAttention (Dao, 2023; Shah et al., 2024) algorithm. FlashAttention reduces memory overhead by breaking the $Q$, $K$, and $V$ matrices into small tiles and fusing matrix multiplication with online softmax computation. When mapping FlashAttention to systolic array based accelerators, the matrix multiplication operations can be performed using the systolic array, while the softmax-related operations must be offloaded to a vector unit. To maximize hardware utilization, matrix multiplication and softmax operations from different iterations are often software pipelined to run in parallel. The vector unit must be able to provide sufficient floating-point operations per second (FLOPs/s) to keep up with the systolic array.
However, vector FLOPs/s are typically much lower than matrix multiplication FLOPs/s, resulting in a significant performance bottleneck. For example, AWS NeuronCore-v2 delivers over $90$ TFLOPs/s for matrix multiplication, but only $2.3$ TFLOPs/s for vector operations (AWS Neuron Documentation, 2024a). This discrepancy becomes especially severe when the systolic array is large, as the matrix multiplication FLOPs/s scale quadratically with the array size, while vector operation throughput remains limited.
Beyond the imbalance between the matrix multiplication and vector FLOPs, another common challenge in implementing FlashAttention on both systolic arrays and GPUs is register file or SRAM port contention. The concurrent execution of matrix and vector computation leads to port contention when both the systolic array and the vector unit attempt to access the same bank of the register file/SRAM, thereby degrading performance. While existing works like Mosaic (Xu et al., 2025) and Virgo (Kim et al., 2025) have aimed to mitigate this contention, they have not fully eliminated it.
The above problems can be addressed if the full FlashAttention algorithm can be executed on a single systolic array, without relying on any external vector unit. In this case, the non-matrix multiplication operations can be performed using the matrix multiplication FLOPs/s of the systolic array, and port contention can be eliminated since all operations are executed on the same array. The key challenges in achieving this are:
  • Efficiently executing non-matrix multiplication operations, such as the exp, rowmax, and rowsum functions, using only the systolic array.
  • Overlapping different operations efficiently to maximize systolic array utilization.
To address these challenges, we propose FSA, an enhanced systolic array architecture that enables the entire FlashAttention algorithm to be executed on a single systolic array, without requiring any external vector unit, based on the following key observations:
  • The exp operation in FlashAttention has a limited range of input values, which allows us to implement it using linear interpolation.
  • The rowmax and rowsum operations can be implemented using systolic array based reduction, which is a well-studied problem. Since the rows are generated in a streaming fashion on the systolic array, the reduction can be performed on-the-fly.
  • After careful study of the dataflow, we observed that the operations of the FlashAttention algorithm can be fully fused and scheduled in a way that allows element-wise operation overlap, enabling the systolic array to perform multiple operations in parallel without waiting for the completion of previous operations.
We further introduce the SystolicAttention scheduling algorithm, which maps FlashAttention onto FSA in a way that enables element-wise operation overlap, significantly improving array utilization.
@@
  • We propose SystolicAttention, a new scheduling algorithm that enables extensive element-wise overlap of all operations in FlashAttention on systolic arrays, while preserving the original floating-point operation order to maintain numerical stability.
  • We design FSA, an enhanced systolic array capable of executing the full FlashAttention algorithm without relying on any additional vector unit. This eliminates port contention entirely and allows ideal performance to be sustained under sufficient memory bandwidth. With our proposed SystolicAttention, FSA achieves 1.77$\times$ and 4.83$\times$ higher FLOPs/s utilization compared to AWS NeuronCore-v2 and TPUv5e, respectively.
  • We fully implement FSA in synthesizable RTL and develop a Python programming interface that allows users to write custom FSA kernels. Both the RTL implementation and the Python software stack are open-sourced at https://github.com/VCA-EPFL/FSA.
This work not only demonstrates the feasibility of executing the entire FlashAttention algorithm on a single systolic array, but also opens up new possibilities for systolic array architectures to handle non-linear functions and other complex operations that were previously considered unsuitable for systolic arrays.
@@ 2. Background and Motivation
In this section, we first briefly introduce systolic arrays and some optimizations that have been proposed based on them. We then present the FlashAttention algorithm and its hardware implementation challenges on systolic array based accelerators.
@@ 2.1. Matrix Multiplication on Systolic Arrays
To perform matrix multiplication on a systolic array, one of the input matrices or output matrix must be preloaded into the array, while the other input matrix is streamed in and the output matrix is streamed out. When an input matrix is preloaded, the array is termed a weight-stationary or input-stationary systolic array. For simplicity, this paper uses the term input-stationary systolic array to refer to both. If the output matrix is preloaded, the array is known as an output-stationary systolic array. Since most deep learning workloads favor input (weight) reuse, most systolic array based accelerators adopt input-stationary architectures.
Figure 1. Computing $C\mathrel{+}=AB$ on an input-stationary systolic array.
@@
An example input-stationary systolic array is shown in Figure 1. It consists of $N\times N$ processing elements (PEs), where each PE can perform a single floating-point multiply-accumulate (MAC) operation per cycle. One of the input matrices is tiled into shape $N\times N$ and preloaded into the array, referred to as the stationary matrix. The other input matrix is tiled into shape $N\times M$ and streamed in from input SRAM, referred to as the moving matrix. The output matrix is thus tiled into shape $N\times M$ and streamed out.
Since the input-stationary array does not retain the output matrix internally, output tiles are stored in a small, dedicated accumulation SRAM located at the bottom of the array. An accumulator unit reads values from this accumulation SRAM, combines the newly computed results from the systolic array with existing values, and writes the updated results back. This near-memory accumulation scheme is widely adopted in both open-source accelerators such as Gemmini (Genc et al., 2022), and commercial products such as AWS NeuronCore (AWS Neuron Documentation, 2024b).
Preloading the stationary matrix requires $N$ cycles at the start of computation. To ensure correctness, each row of the moving matrix must be delayed by 00 to $N-1$ cycles before entering the array, and each column of the output matrix must be delayed by 00 to $N-1$ cycles before being read out. This introduces a total synchronization overhead of $12N-$ 1 cycles. The total computation takes $M+3N-1$ cycles, resulting in an array utilization of $\frac{M}{M+3N-1}$. When $M\gg N$, this preloading and synchronization overhead is effectively amortized over the duration of the computation.
2.2. Systolic Array Optimizations
Several optimizations have been proposed to reduce synchronization overhead in systolic arrays.
Meissa (Asgari et al., 2020) introduces the use of an adder tree for each column in the systolic array to accumulate the results of multiplications. The inputs to the adder tree can be generated in parallel, eliminating the need for synchronization. However, this approach is limited to integer addition, as the adder trees are built using carry-save adders, which are not well-suited for floating-point arithmetic due to the requirements for rounding and exponent alignment.
DiP (Abdelmaksoud et al., 2024) proposes a diagonal-input dataflow to reduce synchronization latency. However, it requires connecting the leftmost PE to the rightmost PE, violating the one hop per cycle communication property that is fundamental to systolic arrays. This alteration increases hardware wiring complexity and latency.
2.3. FlashAttention Algorithm
@@ 22      $O_{i}=diag(old\_l)^{-1}old\_O\in\mathbb{R}^{B_{r}\times d}$;
The FlashAttention algorithm (Dao, 2023; Shah et al., 2024) is shown in Algorithm 1, focusing on the computational steps and omitting memory load/store operations for clarity. The exponential function is implemented as $\exp(x)=2^{x\log_{2}e}$, following the official open-source FlashAttention implementation, since exp2 is typically more efficient than exp on hardware accelerators.
Matrix multiplications are highlighted in red, while non-matrix operations such as reductions and element wise computations are shown in blue. On GPU-like architectures, matrix multiplications are executed on tensor cores, while non-matrix operations are carried out on SIMT cores. On systolic array based accelerators, matrix multiplications are performed using systolic arrays, and the remaining operations are handled by a separate vector unit or scalar core.
To implement FlashAttention on current systolic array based accelerators, the result of the first matrix multiplication must be carried out of the systolic array and into a vector unit to perform the softmax-related operations. The intermediate results must then be transferred back into the systolic array to perform the second matrix multiplication.
Lines 15-16 of Algorithm 1 first compute the local output matrix, which is then accumulated into the global output matrix. The accumulation can be fused with the second matrix multiplication using the accumulator of the systolic array, so it can be treated as a single matrix multiplication operation.
As explained in Section 2.1, additional synchronization and preloading overheads are incurred when moving data in and out of the systolic array. To maximize systolic array utilization, the tile size must be large enough to amortize these overheads. However, a larger tile size also leads to a greater imbalance between the matrix multiplication and vector FLOPs, as well as increased SRAM requirements for storing intermediate results.
Consequently, a significant amount of software tuning is required in addition to software pipelining, and it remains challenging to achieve high performance when implementing FlashAttention on systolic array based accelerators.
3. SystolicAttention Scheduling Algorithm
The goal of our design is to execute the FlashAttention algorithm on a single systolic array, thereby improving the FLOPs/s for the non-matrix multiplication operations and eliminating the need for an external vector unit.
@@ 3. SystolicAttention Scheduling Algorithm
To enable the execution of non-matrix multiplication operations of the FlashAttention algorithm on a single systolic array, we add a row of comparison units at the top of the array to perform the on-the-fly rowmax operation. Additionally, we introduce a Split unit in each PE to implement the exp operation. To allow the rowmax, exp, and rowsum operations to overlap with matrix multiplication, we incorporate an upward data path in the systolic array.
@@ 3.1. FSA Overview
FSA’s three modifications to the standard systolic array are highlighted in red in Figure 2:
A row of comparison units: We add a row of comparison units at the top of the array to compute the maximum value of each column of the systolic array output on the fly. Each comparison unit consists of a floating-point adder and two registers. One register, $old\_m$ holds the maximum value from the previous iteration. The other register, $new\_m$, stores the temporary maximum value of the current iteration. The value of $new\_m$ is updated based on the comparison result, performed by subtracting the two values using the floating-point adder. The difference between the inputs can also be generated using the floating point adder, which is useful to compute Line 9 of Algorithm 1.
A Split unit in each PE: We add a Split unit to each PE to split the input floating-point number into its integer and fractional parts. This is used to implement the exp2 operation in a piecewise linear fashion.
Upward data path: Standard systolic arrays only support propagation of partial sums from matrix multiplication in one direction. We add an additional data path to propagate the partial sums in the opposite direction. This allows computation to begin from either the top or bottom of the array.
@@
Figure 3 illustrates the fused matrix multiplication and rowmax computation, which corresponds to Lines 6–8 of Algorithm 1. The matrix multiplication $S=Q_{i}K^{T}_{j}$ is performed in the systolic array along the upward data path, such that the rows of $S$ become the columns of the systolic array and exit the array from the top.
@@
Consequently, the subtraction operation $N=S-new\_m$ from Line 11 of Algorithm 1 can be performed immediately. This is achieved by streaming constant ones from the left of the array as the multiplicand, and streaming $new\_m$ from the top of the array as the addend.
@@
After the rowmax and subtraction operations, an element-wise exponential function must be applied to the matrix $N$, as shown in Line 12 of Algorithm 1:
@@
A floating-point number $x$ can be split into its integer and fractional parts: $x=x_{i}+x_{f}$, where $x_{i}$ is the integer part and $x_{f}$ is the fractional part (iee, 2019). This leads to the identity:
@@
Since $N=S-\texttt{rowmax}(S)$, all elements in $N$ are less than or equal to zero. This implies $x_{f}\in(-1,0]$, and therefore $2^{x_{f}}\in(0.5,1]$. Given this limited range for $x_{f}$, we find that an 8-piece uniform piecewise linear interpolation is sufficient to achieve a relative error around $10^{-3}$ to $10^{-2}$ in the final FlashAttention results:
@@
To avoid the need for storing the piecewise linear interpolation coefficients in the systolic array, we stream the $slope_{k}$ and $intercept_{k}$ values from the left and top of the array, respectively. We further observe that all intercepts lie in the range $(0.5,1]$, so their exponent can only be 00 or $-1$. The MSBs of their exponents can be used to encode the interpolation index $k$. This enables each PE to update its register based on $k$ without requiring additional control signals.
@@
As shown in Figure 5, the rowsum of $P$ can be computed by streaming constant ones from the left of the array as the multiplicand, and streaming zeros from the top of the array as the addend. The second matrix multiplication, $O=P\cdot V$ can then be performed in the systolic array along the downward data path, starting one cycle after the rowsum operation begins. Once the rowsum $local\_l$ and the output $local\_O$ exit the array from the bottom, they are accumulated with the previous values $old\_l$ and $old\_O$ in the accumulator, and the results update $new\_l$ and $new\_O$ in the accumulation SRAM, respectively.

### v3

@@ Abstract.
Transformer models rely heavily on scaled dot-product attention (SDPA), typically implemented using the FlashAttention algorithm. However, current systolic-array-based accelerators face significant challenges when executing FlashAttention. Systolic arrays achieve high utilization primarily for consecutive and large matrix multiplications, whereas FlashAttention requires frequent interleaving of matrix multiplications and softmax operations.
The frequent data swaps between matrix multiplications on the systolic array and softmax operations on external units result in low array utilization. Moreover, when these computations run concurrently, the softmax stage contends with matrix multiplication for register file and SRAM ports, further degrading performance.
To overcome these limitations, we propose FSA, an enhanced systolic array architecture that enables the FlashAttention algorithm to run entirely within a single systolic array—eliminating the need for external vector units. At the core of FSA is SystolicAttention, a novel scheduling algorithm that maps FlashAttention operations onto systolic arrays with fine-grained, element-wise overlap. This approach significantly improves array utilization while preserving the original floating-point operation order to maintain numerical stability.
[…]
@@ 1. Introduction
The rapid advancement of artificial intelligence, particularly in natural language processing and computer vision, has been largely driven by the success of transformer models (Vaswani et al., 2017; Devlin et al., 2019; Brown et al., 2020), which have revolutionized applications ranging from language translation and text generation (Ouyang et al., 2022) to image recognition (Kolesnikov et al., 2021) and autonomous driving (Carion et al., 2020). However, the computational demands of transformer models are enormous (Strubell et al., 2019), requiring specialized hardware accelerators to achieve practical performance and energy efficiency. As these models continue to grow in size and complexity (Kaplan et al., 2020), the need for efficient hardware acceleration becomes increasingly critical. One of the key operations in transformer models is the scaled dot-product attention (SDPA), which is often implemented using the FlashAttention (Dao, 2023; Shah et al., 2024) algorithm. FlashAttention reduces memory overhead by breaking the $Q$, $K$, and $V$ matrices into small tiles and fusing matrix multiplication with online softmax computation.
Figure 1. Active time percentage of different components in AWS NeuronCore-v2 when running FlashAttention.
[…]
Active Time
@@
Systolic arrays (Kung, 1982) are specialized hardware architectures that perform computation by streaming data through a network of tightly coupled processing elements (PEs). Many systolic array based accelerators with various dataflows (Chen et al., 2017, 2016; Kao et al., 2023; Nayak et al., 2024) have been proposed to accelerate deep learning workloads. In this paper, we focus on multiply-accumulate (MAC) based 2D systolic arrays, which are the most widely used in commercial AI accelerators, such as Google’s TPU (Google Cloud, 2024; Jouppi et al., 2017, 2023) and AWS NeuronCore (AWS Neuron Documentation, 2024b). In these architectures, data moves only one hop per clock cycle, minimizing hardware data movement overhead and enabling high scalability, which is crucial for efficiently executing large matrix multiplications.
@@
When mapping FlashAttention to systolic array based accelerators, the matrix multiplication operations can be executed using the systolic array, while softmax-related operations must be offloaded to external vector and scalar units. To maximize hardware utilization, matrix multiplication and softmax operations from different iterations are often software-pipelined to run in parallel. Thus, the vector and scalar units must provide sufficient floating-point operations per second (FLOPs/s) to keep up with the systolic array. However, non-matrix-multiplication FLOPs/s are typically much lower in volume and more expensive to improve in hardware, resulting in a significant performance bottleneck. As shown in Figure 1, when running FlashAttention on AWS NeuronCore-v2, the systolic array (tensor engine) is only active for about 45% of the time, while the scalar unit remains active for 80% of the time.
Beyond the imbalance between matrix and non-matrix FLOPs/s, the complex data dependencies in FlashAttention also contribute to low systolic array utilization. Systolic arrays achieve high utilization when matrix data can be reused across multiple iterations, improving data locality and reducing the preloading overhead. However, in FlashAttention, the two matrix multiplications are separated by the softmax operation, preventing reuse of intermediate data. The softmax operation also causes frequent data transfers between the systolic array and the vector unit, further increasing preloading overhead and SRAM port contention (Kim et al., 2025; Xu et al., 2025). As shown in our experiments in subsection 6.1, the systolic array achieves less than 25% FLOPs/s utilization, even with 45% active time on AWS NeuronCore-v2, significantly lower than the theoretical peak FLOPs/s of the array.
The above problems can be addressed by executing the full FlashAttention algorithm entirely on the systolic array. In this case, non-matrix-multiplication operations can be computed using the matrix multiplication FLOPs/s available on the array itself. Additionally, data preloading overhead and SRAM port contention are eliminated since all computations occur within the same array. Achieving this, however, requires overcoming two key challenges. First, non-matrix-multiplication operations such as exp, rowmax, and rowsum must be efficiently implemented using only the systolic array. Second, different operations must be overlapped effectively to maximize utilization of the array.
COSA Plus (Wang et al., 2025) proposes mapping FlashAttention to two cooperative systolic arrays, but still requires the softmax operation to be performed by a special function unit (SFU), which may not provide sufficient FLOPs/s to keep up with the systolic arrays. FuseMax (Nayak et al., 2024) pioneered the fusion of FlashAttention into a single systolic array. It maps the rowmax and rowsum operations as spatial reductions and overlaps multiple FlashAttention iterations on the array simultaneously. However, it requires storing intermediate results using 10 registers per PE to support context switching between iterations, leading to additional register overhead.
We propose a new hardware-friendly systolic array architecture, FSA, which introduces minimal overhead to existing commercial systolic arrays and enables the execution of the full FlashAttention algorithm on a single systolic array. This design is guided by the following key insights: (1) the input values to the exp operation in FlashAttention are always less than or equal to zero. By splitting the input into integer and fractional parts, the fractional component falls within the range (1,0](-1,0], making it suitable for approximation via linear interpolation (Koca et al., 2023) reusing the MACs available on the systolic array; (2) the rowmax and rowsum operations can be expressed as reduction operations, which are well-supported in systolic architectures and can be performed on-the-fly as rows are generated in a streaming fashion; and (3) the operations within each FlashAttention iteration exhibit a dataflow pattern that allows fine-grained execution overlap, enabling multiple computations within the same iteration to proceed in parallel without waiting for the completion of prior steps. We also introduce SystolicAttention, a scheduling algorithm that makes the best use of FSA’s capabilities to overlap element-wise operations, significantly improving array utilization.
@@
  • We present FSA, an enhanced systolic array architecture capable of executing the full FlashAttention algorithm without relying on additional vector or scalar units. FSA is implemented in synthesizable RTL, requiring only 12% additional area compared to a standard design. The RTL implementation achieves a clock frequency of 1.5 GHz on a 16 nm process and 70 MHz on the Xilinx Alveo U55C FPGA.
[…]
  • We develop a Python programming interface that enables users to write custom FSA kernels. Both the RTL implementation and the Python software stack are open-sourced at https://github.com/VCA-EPFL/FSA.
@@ 2. Background and Motivation
In this section, we first briefly introduce baseline systolic arrays and how they perform matrix multiplication. We then present the FlashAttention algorithm and discuss the challenges of implementing it on systolic array based accelerators. Finally, we elaborate on the insights and challenges of running the full FlashAttention algorithm entirely within a systolic array.
2.1. Systolic Array Architectures
Systolic arrays are a class of hardware accelerators that perform computations in a pipelined manner, where data flows through an array of processing elements (PEs) in a rhythmic fashion. Each PE performs a simple operation, such as a multiply-accumulate (MAC), and the data flows through the array in a wave-like pattern. Systolic arrays are particularly well-suited for matrix operations, such as matrix multiplication, as they can efficiently exploit data parallelism and locality. The dataflow of a systolic array defines how data moves through the array and the communication pattern between PEs. Common dataflows include weight-stationary, input-stationary, output-stationary, and row-stationary (Chen et al., 2016). Due to the transpose rule of matrix products, weight-stationary and input-stationary dataflows can often be implemented using the same hardware. In this paper, we use the term weight-stationary to refer to both. As most deep learning workloads favor weight reuse, many systolic accelerators adopt the weight-stationary dataflow, such as Google’s TPU (Google Cloud, 2024).
@@ 2.2. Matrix Multiplication on Systolic Arrays
An example weight-stationary systolic array is shown in Figure 2. It consists of $N\times N$ PEs, where each PE performs one floating-point MAC per cycle. One input matrix is tiled into $N\times N$ and preloaded into the array (stationary matrix). The other is tiled into $N\times M$ and streamed in from SRAM (moving matrix). The output is similarly tiled into $N\times M$ and streamed out.
Figure 2. Computing $C\mathrel{+}=AB$ on a weight-stationary systolic array.
@@
Since the weight-stationary array does not retain the output matrix internally, output tiles are stored in a small, dedicated accumulation SRAM located at the bottom of the array. An accumulator unit reads values from this accumulation SRAM, combines the newly computed results from the systolic array with existing values, and writes the updated results back. This near-memory accumulation scheme is widely adopted in both open-source accelerators such as Gemmini (Genc et al., 2022), and commercial products such as AWS NeuronCore (AWS Neuron Documentation, 2024b).
Preloading the stationary matrix requires $N$ cycles at the start of computation. To ensure correctness, each row of the moving matrix must be delayed by 0 to N1N-1 cycles before entering the array, and each column of the output matrix must be delayed by 0 to N1N-1 cycles before being read out. This introduces a total synchronization overhead of 2N12N-1 cycles. The total computation takes M+3N1M+3N-1 cycles, resulting in an array utilization of $\frac{M}{M+3N-1}$. When $M\gg N$, this preloading and synchronization overhead is effectively amortized over the duration of the computation.
@@ 22 $O_{i}=diag(old\_l)^{-1}old\_O\in\mathbb{R}^{B_{r}\times d}$;
The FlashAttention algorithm (Dao, 2023; Shah et al., 2024) is shown in algorithm 1, focusing on the computational steps and omitting memory load/store operations for brevity. Matrix multiplications are highlighted in red, while non-matrix operations such as reductions and element wise computations are shown in blue. The exponential function is implemented as $\exp(x)=2^{x\log_{2}e}$, following the official open-source FlashAttention implementation, since exp2 is typically more efficient than exp on hardware accelerators.
To implement FlashAttention on systolic array based accelerators, the first matrix multiplication result must be moved out to a vector unit or scalar unit for softmax-related operations. Intermediate results must then be sent back into the array for the second matrix multiplication. As discussed in subsection 2.2, these round-trips incur preloading and synchronization overheads. To amortize them, large tile sizes are preferred. However, larger tiles increase the imbalance between matrix and non-matrix FLOPs and raise SRAM demand for intermediate storage. Even with software pipelining to overlap execution, the FLOPs/s imbalance between matrix multiplications and other operations leads to substantial time spent on non-matrix computations (Nayak et al., 2024). Achieving high performance for FlashAttention on systolic arrays remains challenging.
2.4. Motivation
Running the complete FlashAttention algorithm within a single systolic array addresses several critical inefficiencies. By eliminating offloading to external vector or scalar units, all computations can leverage the abundant matrix multiplication FLOPs/s inherently available within the systolic array, significantly improving overall hardware utilization. However, achieving full execution of FlashAttention within the array presents considerable challenges:
  • Efficient Non-MAC Operations: Operations such as exp, rowmax, and rowsum typically rely on external units due to their element-wise or reduction nature. Implementing these efficiently using only MAC-based PEs requires careful algorithmic transformations and architectural extensions.
  • Computation Overlap: To maximize utilization, non-matrix operations must be tightly overlapped with ongoing matrix computations. The complex data dependencies within FlashAttention demand a finely tuned scheduling strategy to ensure that overlapping computations do not stall each other.
  • Dataflow and SRAM Management: The intermediate data produced by the first matrix multiplication in FlashAttention must be effectively reused without excessive SRAM accesses or port contention. Achieving this reuse within the constrained internal storage of the systolic array poses a significant design challenge.
Addressing these challenges requires novel architectural enhancements and sophisticated scheduling algorithms, forming the basis for our proposed FSA architecture and SystolicAttention scheduling algorithm detailed in subsequent sections.
3. FSA Overview
@@
The design principle of FSA is to enable the execution of the full FlashAttention algorithm on a single systolic array with minimal hardware overhead, thereby improving the FLOPs/s for the non-matrix multiplication operations and eliminating the need for an external vector unit. To achieve this, we add a row of comparison units at the top of the array to perform the on-the-fly rowmax operation. Additionally, we introduce a Split unit in each PE to implement the exp operation. To allow the rowmax, exp, and rowsum operations to overlap with matrix multiplication, we incorporate an upward data path in the systolic array.
@@ 3.1. FSA Design Overview
As shown in Figure 3, FSA is a systolic array based accelerator that extends the standard architecture with three key modifications to enable efficient execution of FlashAttention:
  1. (1)
    A row of comparison units at the top of the array to compute the column-wise maximum of the output on the fly.
  2. (2)
    A Split unit in each PE to decompose the input floating-point value into its integer and fractional parts.
  3. (3)
    An upward data path to propagate partial sums in the reverse direction.
The detailed design of these modifications is as follows:
Row of comparison units: To implement lines 11 of algorithm 1, we add a row of comparison units at the top of the array to compute the maximum value of each output column dynamically. Each unit consists of a floating-point adder and two registers. The register old_mold\_m stores the maximum value from the previous iteration, while new_mnew\_m holds the temporary maximum of the current iteration. The value of new_mnew\_m is updated by comparing old_mold\_m and the current input using subtraction through the floating-point adder. This differential computation enables efficient max tracking inline with systolic flow.
Split unit in each PE: Each PE is equipped with a Split unit that separates the input floating-point number into its integer and fractional components. The exp2 operation is applied to the fractional part using piecewise linear interpolation, which is efficiently handled by the PE’s MAC unit.
Upward data path: Standard systolic arrays propagate partial sums in one direction only. We add an additional data path that enables upward propagation, allowing computation to start from either the top or bottom of the array.
These architectural enhancements allow FSA to support operations that are not feasible on a baseline systolic array, thereby enabling efficient standalone execution of FlashAttention, as elaborated in the following sections.
@@
Figure 4 illustrates the fused matrix multiplication and rowmax computation, which corresponds to lines 11 of algorithm 1. The matrix multiplication $S=Q_{i}K^{T}_{j}$ is performed in the systolic array along the upward data path, such that the rows of $S$ become the columns of the systolic array and exit the array from the top.
@@
Consequently, the subtraction operation N=Snew_mN=S-new\_m from algorithm 1 of algorithm 1 can be performed immediately. This is achieved by streaming constant ones from the left of the array as the multiplicand, and streaming new_mnew\_m from the top of the array as the addend.
@@
After the rowmax and subtraction operations, an element-wise exponential function must be applied to the matrix $N$, as shown in algorithm 1 of algorithm 1:
@@
As proposed in (Koca et al., 2023), the input to exp2 can be split into its integer and fractional parts for a fixed-point number, and the fractional part can be approximated using piecewise linear interpolation. We found that this approach is also applicable to floating-point numbers. The piecewise linear interpolation can be efficiently implemented by reusing the MAC units in the systolic array. A floating-point number $x$ can be decomposed into its integer and fractional components: $x=x_{i}+x_{f}$, where $x_{i}$ is the integer part and $x_{f}$ is the fractional part (iee, 2019). This decomposition leads to the identity:
@@
Since $N=S-\texttt{rowmax}(S)$, all elements in $N$ are less than or equal to zero. This implies $x_{f}\in(-1,0]$, and therefore $2^{x_{f}}\in(0.5,1]$. Given this limited range for $x_{f}$, we find that an 8-piece uniform piecewise linear interpolation is sufficient to achieve a relative error around $10^{-2}$ in the final FlashAttention results:
@@
To avoid the need for storing the piecewise linear interpolation coefficients in the systolic array, we stream the $slope_{k}$ and $intercept_{k}$ values from the left and top of the array, respectively. We further observe that all intercepts lie in the range $(0.5,1]$, so their exponent can only be 0 or 1-1. The MSBs of their exponents can be used to encode the interpolation index $k$. This enables each PE to update its register based on $k$ without requiring additional control signals.
@@
As shown in Figure 6, the rowsum of $P$ can be computed by streaming constant ones from the left of the array as the multiplicand, and streaming zeros from the top of the array as the addend. The second matrix multiplication, $O=PV$ can then be performed in the systolic array along the downward data path, starting one cycle after the rowsum operation begins. Once the rowsum local_llocal\_l and the output local_Olocal\_O exit the array from the bottom, they are accumulated with the previous values old_lold\_l and old_Oold\_O in the accumulator, and the results update new_lnew\_l and new_Onew\_O in the accumulation SRAM, respectively.